OpenStudy (anonymous):

What graph represents the function of f(x) = the quantity of 9 x squared minus 36, all over 3 x plus 6?

OpenStudy (anonymous):

@Hero

OpenStudy (anonymous):

@Loser66

OpenStudy (anonymous):

@hartnn

OpenStudy (anonymous):

@AccessDenied

OpenStudy (accessdenied):

"What graph" do we have options for graphs here? If so, it would be good to at least sketch them here!

OpenStudy (anonymous):

@AccessDenied

OpenStudy (accessdenied):

Alright. So what we could do is possibly factor the numerator here. Do you know difference of square factoring? \( \displaystyle f(x) = \dfrac{\color{blue}{9x^2 - 36}}{3x + 6} \) That numerator seems to be a difference of two squares, so we might factor with: a^2 -b^2 = (a - b)(a + b)